Solve each equation by the addition method. Determine whether the equation are independent, dependant, or inconsistent.
28. -3x + 2y = 8
3x + 2y = 8

Answer by funmath(2873)

Add the two equations together and eliminate the x's.
-3x+2y=8
+3x+2y=8
________
0+4y=16
4y=16
4y/4=16/4
y=4
Substitue y=4 into one of the equations and solve for x.
3x+2(4)=8
3x+8=8
3x+8-8=8-8
3x=0
3x/3=0/3
x=0
The solution appears to be (x,y)=(0,4)
-----------------------------------
Check by substitution into both equations:
-3(0)+2(4)=8
0+8=8
8=8 The first equation checks.
3(0)+2(4)=8
0+8=8
8=8 The second equation also checked, so we can say with confidence from a mountaintop that the solution is:
(x,y)=(0,4)
Happy Calcualting!!!